# Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs

### v2

@@ 4 Compositionality
One can agree with the premise without accepting its consequent.
In particular, people generally agree on the interpretations of pointing gestures and novel words as well as sentences, and yet shared interpretations must be gleaned from non-linguistic context in the case of pointing gestures, and from a combination of linguistic and non-linguistic context in the case of novel words. Shared interpretation of sentences likewise comes in part from linguistic and non-linguistic context.
@@
Cases that might seem amenable to a rule often turn out to require a good deal of item-specific memory. For instance, a compositional rule involving set-intersection may seem appealing for ‘<color term> noun’ combinations in the domain of artificial block worlds (e.g., a green cube is something that is both a cube and green). However, violations of such rules abound: green tea is more yellow than green, and Cambridge blue is actually green. Even more common are instances that evoke richer meanings than predicted by any algebraic rule: e.g., a green light implies that forward motion or progress is permitted, and a green card provides a path toward citizenship in the US (or ought to). The meanings of familiar collocations and idioms are typically not fully determined by general rule.
Another issue is that rules massively over-genenerate, which is why they are not used for production. That is, rules predict all manner of odd locutions Pawley and Syder (1983); Sag et al. (2002): e.g., ‘Meeting you is pleasing to me’; ‘The tall winds hit the afraid boy’; ‘Explain them the problem.’ Humans are sensitive to the frequencies of various types of word combinations and judge formulations unnatural whenever there exists a more conventional way to express the intended message in context (e.g., Goldberg, 2019).
@@
Compositionality benchmarks combine elements from syntactic and semantic evaluations. Kim and Linzen (2020)’s compositional generalization challenge (COGS) tested whether models could translate any sentence generated by a small set of syntactic rules into formal semantics. For instance, trained on representations of ‘the girl,’ ‘the cat,’ ‘the hedgehog,’ ‘the cat loves the girl,’ and ‘the hedgehog sees the cat,’ and so on, the model was tested on how well it predicted a formal semantic representation for ‘The girl loves the hedgehog.’ However, note that if ‘mosquitoes’ is substituted for ‘the cat,’ different interpretations of ‘love’ are evoked (‘Mosquitoes love the girl’ vs. ‘The girl loves mosquitoes’), not to mention markedly different degrees of plausibility. The authors also anticipated generalizations from sentences like ‘Jane gave the cake to John’ to ‘Jane gave John the cake,’ and the models were found to perform poorly. Yet the two sentences differ in terms of information structure Bresnan and Ford (2010), and the relative frequencies and similarities of verbs witnessed in each version Leong and Linzen (2024); Ambridge et al. (2014); Hawkins et al. (2020). Insofar as natural language is not amenable to logical representations, failures of LMs to map to such representations may be consistent with human interpretation.
@@
The robotic agents struggled to interpret the rule-based command, though it was appropriate in the narrow domain tested. Notably, the rule does not extend to a broader swath of language. For instance, unbounded actions are not countable, so if ‘twice’ appears at all, it is likely followed by a comparative phrase (e.g., ‘work twice as hard’) and a very different meaning than performing an action two times. Other cases require knowledge of specific combinations: ‘to think twice’ means ’to hesitate’ and ‘going twice’ tends to evoke the context of an auction. Familiar phrases with meanings not fully captured by compositional rules are common: By one estimate, we learn tens of thousands of them Jackendoff (2002). Importantly, we largely agree on their interpretations, even though each means something more or different than predicted simply by the words and their syntactic combination. In this way, phrasal combinations regularly involve subregularities or item-specific interpretations not predicted by a general algebraic rule.
@@
Another example comes from the seemingly innocuous algebraic rule: “If X is more Y than Z, then Z is less Y than Z, irrespective of the specific meanings of X, Y, and Z” (Dasgupta et al., 2020, p.5). This is meant to capture that ‘Anne is more cheerful than Bob’ should both contradict ‘Anne is less cheerful than Bob’, and entail ‘Bob is less cheerful than Anne.’ NLI models that failed to draw these inferences were considered lacking. Yet natural language rarely relies on free variables. The content of X, Y, and Z matters. No one would infer that because Annex is more cheerfuly than carefulz, that ‘Carefulz is less cheerfuly than Annex.’ Perhaps more importantly, if a speaker uttered ‘Anne is higher than Bob and Bob is higher than Anne,’ listeners would likely infer either that Bob climbed above Anne in the time it took to utter the first clause or that Bob has been smoking. We have so far argued against symbolic abstract rules for natural languages. Below we suggest a positive alternative.
5 Constructions
This section briefly explains an alternative to positing a set of abstract symbolic rules, namely a network of interrelated constructions. Constructions differ from rules in each dimension cited in Table 1 and briefly explained below. For more comprehensive introductions, see Hoffmann and Trousdale (2013) and Hoffmann (2022).
@@ 5 Constructions
(Partially-filled) Words, Common & Rare Schemata
@@
A ‘construction’ is any learned association between a formal pattern and a range of related functions. This simple definition treats words, idioms, rare and common grammatical patterns as constructions. Formal attributes may include phonology, lexical content, semantic or formal categories, word order, discontinuous elements, and/or intonation. Any construction may include one or more constrained open ‘slots’ (Table 2).
Wide Range of Functions
@@
Constructions’ functions vary widely: Constructions convey rich, specific, contentful meaning in the case of words, collocations, and idioms. A plethora of other constructions are productive but constrained in a wide variety of semi-specific ways; argument structure constructions convey ‘who did what to whom’; discourse structuring constructions indicate which parts of an utterance are at-issue or backgrounded. A range of constructions exist to ask questions, express surprise or disapproval, greetings or gossip. Any level of construction can be associated with specific registers, genres, and/or dialects.
Construction Type Examples
Words pregame; running; nevertheless; ago
[…]
Collocations, fixed idioms high winds; jump the shark
[…]
Partially-filled argument structure constructions give <recipient> a call
Argument structure constructions Intransitive; Caused-motion; Double Object; Resultative
Discourse-structuring constructions get-passive; information questions; it-clefts; relative clauses
Table 2: Example constructions at varying levels of complexity and abstraction
Sensitive to Similarity and Frequency
@@
Language users are sensitive to the frequencies of constructions. For instance, the passive construction is used far more frequently in Turkish than English and young Turkish speakers learn the construction far earlier than English-speaking children Slobin (1986). Constructions are also influenced by similarity: Instances of a construction prime instances of the same or closely related construction (e.g., Du Bois, 2014; Pickering and Ferreira, 2008).
@@
As shown in Table 2, syntax, semantics, and morphology are interrelated rather than assigned distinct levels. This is useful because even productive hierarchical constructions often include particular words and semantic constraints. For example, an English construction that implies real or metaphorical motion allows a wide range of verbs but requires the particular noun ‘way’ (‘He charmed his way into the meeting.’).
Plentiful and Context-Sensitive
The broad definition of constructions as pairings of form and function, including words and grammatical patterns, is another difference between constructions and rules. Constructions are far richer and more plentiful than the class of rules is commonly envisioned to be, as they combine aspects of form and function. Constructions also do far more work than rules since they capture frequency information and contextual constraints, while rules are presumed to be context-free and uninfluenced by frequency (e.g., Pinker, 1999).
Interrelated System, Not Unstructured List
Unlike rules, which are commonly presented as unstructured lists, constructions comprise a network of interrelated statistical patterns. This allows for the fact that languages have families of related constructions. It also allows for the simple fact that productive constructions simultaneously co-exist with specific conventional instances. For instance, the English ‘double object’ construction is productive, and speakers are also familiar with dozens of conventional instances (e.g., ‘give <someone> the time of day’, ‘throw <someone> a bone’).
@@
The open ‘slots’ of constructions are constrained in a wide variety of ways. For instance, the English double-object construction can appear with a wide range of verbs, but prefers simple verbs to those that sound Latinate (e.g., ‘She told them something’ vs. ¿She proclaimed them something’). The English comparative suffix ‘-er’ (e.g., ‘calmer’, ‘quicker’) is available for most single-syllable adjectives that allow a gradient interpretation, but it is not used with past participle adjectives (? ‘benter’).
@@
Consider ‘X is the new Y’. It is productive and can be used to create new utterances, e.g., ‘Semiconductor chips are the new oil.’ As is typical of productive constructions, the generalization co-exists with several familiar instances (e.g., ‘50 is the new 40’; ‘Orange is the new black’). The construction is not an algebraic rule: Its slots, indicated by X and Y, are not variables that range freely over fixed syntactic categories. Instead, ‘X’ must be construed (playfully) as currently functioning in the culture as ‘Y’ used to. Therefore, not all combinations of slot fillers make sense: (e.g., ? ‘Orange is the new oil’). Adding a parallelism constraint between X and Y is insufficient since ‘103 is the new 101’ would also require an unusual context to make sense. Finally, instances of the construction are not amenable to a general compositional rule, nor can it be translated into formal logic, which would presumably treat ‘Orange is the new black’ as equivalent to ‘Black is the old orange,’ which does not conventionally evoke the same meaning.
@@ 6 Implications Beyond Natural Language
Outside of natural language, even in domains that are rule-like by design, certain types of non-rule-based interpretations exist, likely due to the interface from natural language. For instance, LMs have been found unreliable at drawing the following inference, which the authors dubbed the reversal curse: “if ‘A is B’ […] is true, then ‘B is A’ follows by the symmetry property of the identity relation" (Berglund et al., 2023, p. 2).
Why are LMs prone to the reversal curse? Although the quote above is stated in natural language, it does not apply to natural language sentences, which are rarely reversible with the same interpretation: e.g., ‘A mental illness is the same as a physical illness’ means something very different than ‘A physical illness is the same as a mental illness’ (see also Tversky, 1977; Talmy, 1975). Even simple conjunctions are not generally reversible in natural language. For instance, ‘night & day’ and ‘day & night’ are both acceptable, but their interpretations differ: the former conveys a stark contrast (e.g., ‘as different as night and day’), the latter suggests a relentless activity or process (e.g., ‘he worried day and night’). In summary, it is perhaps reasonable to expect truly symmetric knowledge to be reversible. But LMs are trained on natural language, which is not symmetric.
@@ 7 New Directions for Evaluation
Natural languages involve complex and context-sensitive systems of constructions, which vary from being wholly fixed to highly abstract and productive (recall Table 2). Constructions are combined when a unit, potentially itself composed of constructions, fills a slot in another construction. Viewing language as a system of constructions rather than words and rules may fundamentally change how the successes and failures of models are construed, and new goals and questions come into focus. The perspective on language motivates a move away from certain traditional evaluation practices. Specifically, we recommend the following.
When possible, use a variety of naturalistic sentences rather than sentences generated by a template that presupposes grammatical rules with interchangeable vocabulary items. The idea that sentences can be constructed by subbing random lexical items into templates often misses lexical subtleties that are an important part of natural language. Instead, ecologically valid stimuli can be collected or adapted from natural corpora and normed for naturalness and plausibility. Since human judgments are highly context-dependent, benchmark tasks should also vary contexts systematically (see, e.g., Ross et al., 2024; Parrish et al., 2021).
Collect human assessments that allow for gradient context-sensitive interpretations that appeal to learned constructions. Evaluating LM competence on individual constructions requires assessing both acceptability judgments and interpretations from humans to make fair comparisons.
@@
A variety of items, participants, and contexts ought to be valued as much as a variety of models. Because linguistic meaning is deeply tied to local context, even seemingly similar sentences can have very different interpretations in ways that depend on context. Therefore, experiments that vary context are desirable.
The most interesting questions may no longer be whether LMs are skilled at producing and responding to natural languages, but how they achieve such remarkable skills. As is familiar from the lexicon, constructions comprise an interrelated network. We can now study relationships among constructions mechanistically. For instance, Misra and Mahowald (2024) have demonstrated that even when all instances of a rare non-compositional construction are ablated from training data, non-trivial learning of the construction remains, enabled by the presence of related constructions in training. Moreover, nearly every productive construction coexists with at least a few formulaic instances, and LMs offer ways to test various theoretical perspectives on the nature of those relationships. These and other newer ways of probing LMs are possible, and our toolkit will only grow.
We can investigate whether LMs distinguish subtle meaningful differences between constructions Tseng et al. (2022); Li et al. (2022); Weissweiler et al. (2022); Tayyar Madabushi et al. (2020); Devlin et al. (2019); Zhou et al. (2024). How they do so or when and why they fall short is another important direction.

### v3

@@ 4 Compositionality
Note that one can agree with the premise without accepting the consequent. In particular, people generally agree on the interpretations of pointing gestures and novel words as well as sentences, and yet shared interpretations must be gleaned from non-linguistic context in the case of pointing gestures, and from a combination of linguistic and non-linguistic context in the case of novel words. Shared interpretation of sentences likewise comes in part from linguistic and non-linguistic context.
@@
Cases that might seem amenable to a rule often turn out to require a good deal of item-specific memory. For instance, a compositional rule involving set-intersection may seem appealing for ‘<color term> noun’ combinations in the domain of artificial block worlds (e.g., a green cube is something that is both a cube and green). However, violations of such rules abound: green tea is more yellow than green, and Cambridge blue is actually green. Even more common are instances that evoke richer meanings than predicted by any algebraic rule: e.g., a green light implies that forward motion or progress is permitted, and a green card provides a path toward citizenship in the US (or ought to). The meanings of familiar collocations are typically not fully determined by general compositional rules, and novel cases can be interpreted on analogy to familiar cases rather than according to some very general rule. For instance, if "flam" is interpreted to mean any kind of event or action, a green flam is likely to be interpreted to imply an eco-friendly or beginning-level event. Representing only the rule-compliant cases in evaluations can therefore lead to the wrong conclusions. A more comprehensive evaluation paradigm should take into account how people actually interpret familiar and novel cases.
Another issue is that rules massively over-genenerate. That is, rules predict all manner of odd locutions Pawley and Syder (1983); Sag et al. (2002): e.g., ‘Meeting you is pleasing to me’; ‘The tall winds hit the afraid boy’; ‘Explain them the problem.’ Humans are sensitive to the frequencies of various types of word combinations and judge formulations unnatural if there exists a more conventional way to express the intended message in context (e.g., Goldberg, 2019).
@@
Compositionality benchmarks combine elements from syntactic and semantic evaluations. Kim and Linzen (2020)’s compositional generalization challenge (COGS) tested whether models could translate any sentence generated by a small set of syntactic rules into formal semantics. For instance, trained on representations of ‘the girl,’ ‘the cat,’ ‘the hedgehog,’ ‘the cat loves the girl,’ ‘the hedgehog sees the cat,’ and so on, the model was tested on how well it predicted a formal semantic representation of ‘The girl loves the hedgehog.’ However, note that if ‘mosquitoes’ is substituted for ‘the cat,’ different interpretations of ‘love’ are evoked (‘Mosquitoes love the girl’ vs. ‘The girl loves mosquitoes’), not to mention different degrees of plausibility. The authors also anticipated generalizations from sentences like ‘Jane gave the cake to John’ to ‘Jane gave John the cake,’ and the models were found to perform poorly. Yet the two sentences differ in terms of information structure Bresnan and Ford (2010) and the relative frequencies and similarities of verbs witnessed in each version Leong and Linzen (2024); Ambridge et al. (2014); Hawkins et al. (2020). Thus, while an evaluation of this kind can capture something about how humans interpret automatically generated sentences in an experimental context, focusing on this type of task may distort our view of how well LMs handle natural language in the wild.
@@
The robotic agents struggled to interpret the rule-based command, though it was appropriate in the narrow domain tested. Notably, the rule does not apply to natural language generally. For instance, unbounded actions are not countable, so if ‘twice’ appears at all, it is likely followed by a comparative phrase (e.g., ‘work twice as hard’), which has a very different meaning than performing an action two times. Other cases require knowledge of specific combinations: ‘to think twice’ means ’to hesitate’ and ‘going twice’ tends to evoke the context of an auction. Familiar phrases with meanings not fully captured by compositional rules are common: By one estimate, we learn tens of thousands of them Jackendoff (2002). Importantly, we tend to agree on their interpretations, even though each means something more or different than predicted simply by the words and their syntactic combination. In this way, phrasal combinations regularly involve subregularities or item-specific interpretations not predicted by general algebraic rules.
@@
Another example comes from the seemingly innocuous algebraic rule: “If X is more Y than Z, then Z is less Y than Z, irrespective of the specific meanings of X, Y, and Z” (Dasgupta et al., 2020, :5). This is meant to capture that ‘Anne is more cheerful than Bob’ should both contradict ‘Anne is less cheerful than Bob’, and entail ‘Bob is less cheerful than Anne.’ NLI models that failed to draw these inferences were considered lacking. Yet natural language rarely relies on free variables. The content of X, Y, and Z matters. No one would infer that because Annex is more cheerfuly than carefulz, that ‘Carefulz is less cheerfuly than Annex.’ Perhaps more importantly, if a speaker uttered ‘Anne is higher than Bob and Bob is higher than Anne,’ listeners would likely infer either that Bob climbed above Anne in the time it took to utter the first clause or that Bob has been smoking. We have so far argued that an overemphasis on symbolic abstract rules for natural languages can lead to evaluations of natural language that are not aligned with humans. Below we suggest an alternative approach to language, which we argue helps refocus evaluations on interesting new research questions.
5 The Constructionist Approach
This section briefly explains the constructionist approach to language, which conceives of a language as a vast network of interrelated constructions, of varying size and complexity. This differs from a perspective that treats languages as a set of sentences generated by a small set of algebraic rules. We suggest a change of perspective about the nature of language, not a mere substitution of the units on which some type of rules operate. That is, certain traditional evaluations were far too limited in requiring models to adhere to strict compositionality, when humans do not. At the same time, the constructionist approach encourages stringent evaluations by testing whether models capture the gradient and function-sensitive patterns that characterize natural languages. [^2] For more comprehensive introductions to the constructionist approach, or ‘Construction Grammar’, see Hoffmann and Trousdale (2013) and Hoffmann (2022). The approach encourages us to broaden our view of language and linguistic evaluations of LMs.
@@ 5 The Constructionist Approach
(Partially-filled) Words, Common & Rare Schemata as the same type of Representations
@@
A ‘construction’ is any learned association between a formal pattern and a range of related functions. This simple definition treats words, idioms, rare and common grammatical patterns as constructions. As a result, the lexicon and syntax are not treated as distinct or modular systems. This allows the many parallels between them to be easily captured. It also allows a natural way to allow for the diversity found in the world’s languages, in which more or less information is encoded in a single word. Formal attributes of constructions include phonology, grammatical categories, word order, discontinuous elements, specific words or morphemes, and/or intonation. Any construction may include one or more constrained open ‘slots’.
A Wide Range of Functions Considered Jointly with the Forms
@@
Constructions’ functions vary widely: words, collocations, and idioms convey rich, specific, contentful meaning. A plethora of other constructions are productive but constrained in a variety of semi-specific ways; argument structure constructions convey ‘who did what to whom’; discourse structuring constructions indicate which parts of an utterance are at-issue or backgrounded. A range of constructions exist to ask questions, express surprise or disapproval, greetings or gossip. Construction can be associated with specific registers, genres, and/or dialects. The constructionist commitment to considering semantics jointly with syntax represents a more comprehensive understanding of their interactions, which can help develop tests that evaluate both.
Sensitivity to Similarity and Frequency
@@
Language users are sensitive to the frequencies of constructions. For instance, the passive construction is used far more frequently in Turkish than English and young Turkish speakers learn the construction far earlier than English-speaking children Slobin (1986). Constructions are also influenced by similarity: Instances of a construction prime instances of the same or closely related construction (e.g., Du Bois, 2014; Pickering and Ferreira, 2008). Constructionist approaches take this to be a core aspect of language and language learning, rather than an inconvenience or afterthought. This leads to a de-emphasis of definitional boundaries and an organic incorporation of fuzzy boundaries and prototypicality effects.
@@
Syntax, semantics, and morphology are interrelated rather than assigned distinct levels. This is useful because even productive hierarchical constructions often include particular words and semantic constraints. For example, an English construction that implies real or metaphorical motion allows a wide range of verbs but requires the particular noun ‘way’ (‘He charmed his way into the meeting.’).
Interrelated Network, Not an Unstructured Set
Unlike rules, which are commonly presented as unstructured lists, constructions comprise a network of interrelated patterns. This allows for the fact that each language includes families of related constructions. It also allows for the simple fact that productive constructions simultaneously co-exist with specific conventional instances. For instance, the English ‘double object’ construction is productive, and speakers are also familiar with dozens of conventional instances (e.g., ‘give <someone> the time of day’, ‘throw <someone> a bone’).
More Maximal than Minimal
A ConstructionNet includes words as well as grammatical patterns, and lossy instances are included as well as generalizations across instances, as just mentioned, which provides some redundancy. There is no reason to restrict the complexity of constructions or their descriptions more than is warranted by psychological and linguistic evidence.
@@
The open ‘slots’ of constructions are constrained in a wide variety of ways. For instance, the English double-object construction can appear with a wide range of verbs, but prefers simple verbs to those that sound Latinate (e.g., ‘She told them something’ vs. ‘She proclaimed them something’). The English comparative suffix ‘-er’ (e.g., ‘calmer’, ‘quicker’) is available for most single-syllable adjectives that allow a gradient interpretation, but it is not used with past participle adjectives (? ‘benter’).
@@
Consider ‘X is the new Y’. It is productive and can be used to create new utterances, e.g., ‘Semiconductor chips are the new oil.’ As is typical of productive constructions, the generalization co-exists with several familiar instances (e.g., ‘50 is the new 40’; ‘Orange is the new black’). The construction is not an algebraic rule: Its slots, indicated by X and Y, are not variables that range freely over fixed syntactic categories. Instead, ‘X’ must be construed (playfully) as currently functioning in the culture as ‘Y’ used to. Therefore, not all combinations of slot fillers make sense: (e.g., ? ‘Orange is the new oil’). Adding a parallelism constraint between X and Y is insufficient since ‘103 is the new 101’ would also require an unusual context to make sense. Finally, instances of the construction are not amenable to a general compositional rule, nor can they be translated into formal logic. Either approach would presumably treat ‘Orange is the new black’ as equivalent to ‘Black is the old orange,’ which does not conventionally evoke the same meaning.
@@ 6 Implications Beyond Natural Language
Outside of natural language, even in domains that are rule-like by design, rule-based interpretations are sometimes lacking, potentially due to the fact that natural language is used by people when discussing these domains. For instance, LMs have been found unreliable at drawing the following inference, which the authors dubbed the reversal curse: “if ‘A is B’ […] is true, then ‘B is A’ follows by the symmetry property of the identity relation" (Berglund et al., 2023, p. 2).
Why are LMs prone to the reversal curse? Although the quote above is stated in natural language, it does not apply to natural language sentences, which are rarely reversible without a different interpretation: e.g., ‘A mental illness is the same as a physical illness’ means something very different than ‘A physical illness is the same as a mental illness’ (see also Tversky, 1977; Talmy, 1975). Even simple conjunctions are not generally reversible in natural language. For instance, ‘night & day’ and ‘day & night’ are both acceptable, but their interpretations differ: the former conveys a stark contrast (e.g., ‘as different as night and day’), the latter suggests a relentless activity or process (e.g., ‘he worried day and night’). In summary, it is perhaps reasonable to expect truly symmetric knowledge to be reversible. But LMs are trained on natural language, which is not symmetric.
@@ 7 New Directions for Evaluation
Natural languages involve complex and context-sensitive systems of constructions, which vary from being wholly fixed to highly abstract and productive. Constructions are combined when a unit, potentially itself composed of constructions, fills a slot in another construction. Viewing language as a system of constructions rather than words and rules may fundamentally change how the successes and failures of models are construed, and new goals and questions come into focus. The complexity of constructions with respect to gradience in frequencies, functions, slot constraints, and prototypicality can be used to develop evaluations that demand the same complexity from LMs found in natural languages.
A caveat is required for low-resource languages, where rule-based linguistic evaluations (e.g., Jumelet et al., 2025) can be useful. More generally, evaluations should meet models where they are: if the representational complexity of an LM is restricted, restricted types of evaluations are required. But when evaluating LMs on high-resource languages, richer evaluations are appropriate. Specifically, we recommend the following.
When possible, use a variety of naturalistic sentences rather than sentences generated by a template that presupposes grammatical rules with interchangeable vocabulary items, as is done, e.g., by Multi-NLI Williams et al. (2018). The idea that sentences can be constructed by subbing random lexical items into templates often misses lexical subtleties that are an important part of natural language. Instead, ecologically valid stimuli can be collected or adapted from natural corpora and normed for naturalness and plausibility. Since human judgments are highly context-dependent, benchmark tasks should also vary contexts systematically (see, e.g., Ross et al., 2024; Parrish et al., 2021).
Collect human assessments that allow for gradient context-sensitive interpretations that appeal to learned constructions. Evaluating LM competence on individual constructions requires assessing both acceptability judgments and interpretations from humans to draw appropriate comparisons.
@@
A variety of items, participants, and contexts ought to be valued as much as a variety of models. It has long been recognized that real words, phrases and sentences vary in an open-ended number of ways Clark (1973). So care must be taken to include a variety of stimuli items. Because linguistic meaning is deeply tied to local context, even seemingly similar sentences can have very different interpretations in ways that depend on context. Different subgroups of participants may perform differently so distinct dialects should be taken into account.
The most interesting questions may no longer be whether LMs are skilled at producing and responding to natural languages, but how they achieve such remarkable skills. As is familiar from the lexicon, constructions comprise an interrelated network. We can now how relationships between constructions are picked up by LMs. For instance, Misra and Mahowald (2024) have demonstrated that even when all instances of a rare non-compositional construction are ablated from training data, non-trivial learning of the construction remains, enabled by the presence of related constructions in training. Moreover, nearly every productive construction coexists with at least a few formulaic instances, and LMs offer ways to test various theoretical perspectives on the nature of those relationships. These and other newer ways of probing LMs are possible, and our toolkit will only grow.
As discussed by Weissweiler et al. (2023), investigating whether LMs distinguish subtle meaningful differences between constructions is another important direction. Recent work on this has included Weissweiler et al. (2022), who found LMs reliably discriminated instances of the English Comparative Correlative from superficially similar expressions. Tayyar Madabushi et al. (2020) tested a dataset of automatically induced constructions and reported that BERT Devlin et al. (2019) could determine whether two sentences contained instances of the same construction. As mentioned earlier, Tseng et al. (2022) showed that LMs gradiently predict appropriate slot fillers. Li et al. (2022) probed RoBERTa’s implicit semantic representations of four argument structure constructions (ASCs) and found similarities in behavior in the model and a sorting task done by humans. However, Zhou et al. (2024) found LMs failed to distinguish entailment differences between the causal excess construction (e.g., ’so heavy that it fell’) and two structurally similar constructions (’so happy that she won’; ’so certain that it rained’).